# A neural network for estimating compact binary coalescence parameters of gravitational-wave events in real time

### v1

@@ Abstract
Low-latency pipelines analyzing gravitational waves from compact binary coalescence events rely on matched filter techniques. Limitations in template banks and waveform modeling, as well as non-stationary detector noise cause errors in signal parameter recovery, especially for events with high chirp masses. We present a quantile regression neural network model that provides dynamic bounds on key parameters such as chirp mass, mass ratio, and total mass. We test the model on various synthetic datasets and real events from LIGO-Virgo-KAGRA gravitational-wave transient catalogs. We find that the model accuracy is consistently over 90% across all the datasets. We explore the possibility of employing the neural network bounds as priors in online parameter estimation. We find that they reduce by 9% the number of likelihood evaluations. This approach may shorten parameter estimation run times without affecting sky localizations.
@@
- Laser Interferometer GW Observatory
@@ I Introduction
Over the course of three observing runs, LIGO, Virgo, and KAGRA (LVK) have confirmed about compact binary coalescence (CBC) detections of binary neutron stars, binary black holes and neutron star-black holes Abbott et al. (2019, 2021a, 2021b). The fourth observing run (O4) 111https://observing.docs.ligo.org/plan has seen more than 200 public alerts of gravitational-wave (GW) triggers 222https://gracedb.ligo.org/superevents/public/O4/.
@@ II Data and model architecture
The second dataset is taken from the LVK third observing run (O3) replay Mock Data Challenge (MDC) Sharma Chaudhary et al. (2024). It consists of simulated CBC signals that were analyzed by online LVK search pipelines before the O4 in order to evaluate their performance. We utilize this dataset solely for testing purposes due to the limited number of simulated signals that the pipelines successfully recover. We also utilize a subset of the O3 replay MDC data for our PE analysis in subsection VII.
@@
As the chirp mass of the signal increases, the error in its pipeline recovered value also increases Ewing et al. (2023). Therefore, following a binning approach adopted by the online PE algorithm bilby-pipe 1.3.0
, we split the data into four different bins according to the value of the chirp mass. Table 1 lists the lower and upper bounds of these bins as well as their sizes. In order to train and test our method, we further split the datasets into () train, validation, and test sets.
@@
In the case of a single NN model predicting multiple quantiles , where , the loss function can be generalized to
Our NN QR configuration is illustrated in Fig. 1 and described below following the PyTorch Ansel et al. (2024) library nomenclature Foundation (2025). The architecture consists of a fully connected NN with two (24, 12) hidden layers interleaved with dropout and LeakyReLU
activations. The input is pre-processed by a LayerNorm
layer. The output of the hidden layers is sorted in ascending order by a SoftSort
@@
. This configuration prevents possible inversions of quantile values in the output Blondel et al. (2020). The latter is a set of unconstrained real numbers that is conditioned according to the specific training parameter: chirp mass, , mass ratio, , and total component mass . For the mass ratio we apply a sigmoid function to constrain the output of the sorting layer to the physical range . In the case of chirp mass and total component mass, we take the exponential of the sorting layer output and multiply it by the corresponding recovered parameters. For example, in the case of the chirp mass we have , where the prime indicates the output of the sorted layer.
We utilize the AdamW
optimizer, the CosineAnnealingWarmRestarts learning rate scheduler, and initial learning rates equal to () for and (). We train with the stochastic gradient descent and loop through the dataset in 100 batches of samples (epochs). The model weights are selected across epochs to minimize the loss on the validation set. The optimal training hyperparameters are selected by testing multiple configurations of the optimizer, learning rate, and the number of nodes in the hidden layers. The final choice is based on the configuration that minimizes the validation loss.
@@
Figure 2 displays the loss curves for the twelve independent models (four chirp mass bins three parameters). As the number of the training epoch increases, both training and validation losses decrease, becoming essentially constant after 50 epochs. This is due to the annealing of the learning rate.
@@
To assess the performance of our models we define two metrics, accuracy and interval width. For each quantile with target accuracy we build the confidence interval , where is any of the training parameters. If the true (injected) value of the parameter lies within the interval, we call it accurate. The fraction of accurate intervals across the dataset defines the accuracy of the models. The interval width is defined as . An optimal model produces intervals with maximum accuracy and smallest width compatible with the uncertainties of the recovered parameter.
@@ III O2 results
Figure 3 shows the performance of the models on the O2 testing dataset. The dashed diagonal lines indicate ideal performance. The model for the mass ratio slightly underperforms for the A and C bins. This could be an effect of the final sigmoid layer for the model. The sigmoid function only asymptotically reaches and . Therefore, the NN output may be less accurate for equal or extreme mass ratio events.
@@ IV O3 MDC results
0.954 | 70.291 |
Figure 5 and Table 3 show the results of the NN models trained on O2 data on the O3 MDC set. Since the parameter space of the O2 and the O3 MDC datasets are different, we evaluate the models on a reduced O3 dataset (see Sec. II.) The different trends of the parameters in the bins wrt O2 are likely due to the inclusion of multiple pipelines in the O3 dataset, while training on the output of a single pipeline (GstLAL) in the O2 data. Changes in pipelines and their template banks leading up to O4 also affect injection recovery across different datasets. In spite of this, Fig. 5 and table 3 confirm that the NN models perform similarly on the O3 set as they do on O2.
@@ V Performance on a synthetic dataset
In order to assess the NN performance on the full parameter space, we create a dataset of synthetic injections. We choose a log-uniform distribution for between 0.5 and 55 and a uniform distribution for between 0.05 and 1. We set the spin components to zero and fix the SNR to 10, 15, 20, and 25. Figure 7 shows the NN interval width with 96% target accuracy. Overall, in each bin the interval width decreases as increases. The width also generally increases with at fixed . These trends can also be seen in Fig. 8 which shows the widths of the chirp mass intervals for different values of the SNR. The median widths decrease with increasing SNR and other input parameters held fixed. This is more evident for bins B, C, and D.
@@ VII NN model outputs in low-latency PE
We produce full posteriors with bilby 2.2.1
and bilby-pipe 1.3.0
and uniform mass priors with intervals computed by the NN. We also perform PE on the same set with the default low-latency bilby listed in Table 5. We employ the relative binning likelihood Krishna et al. (2023) since the low-latency ROQ method is incompatible with the dynamical chirp mass intervals provided by the NN.
@@
Figure 9 compares the number of likelihood evaluations for Bilby’s default configuration and with the NN priors. The NN priors show a % median reduction in the number of likelihood evaluations.

### v2

@@ Abstract
Low-latency pipelines analyzing gravitational waves from compact binary coalescence events rely on matched filter techniques. Limitations in template banks and waveform modeling, as well as non-stationary detector noise cause errors in signal parameter recovery, especially for events with high chirp masses. We present a quantile regression neural network model that provides dynamic bounds on key parameters such as chirp mass, mass ratio, and total mass. We test the model on various synthetic datasets and real events from the LIGO-Virgo-KAGRA gravitational-wave transient GTWC-3 catalog. We find that the model accuracy is consistently over 90% across all the datasets. We explore the possibility of employing the neural network bounds as priors in online parameter estimation. We find that they reduce by 9% the number of likelihood evaluations. This approach may shorten parameter estimation run times without affecting sky localizations.
@@
- Laser Interferometer Gravitational-Wave Observatory
@@ I Introduction
Over the course of three observing runs, the LIGO, Virgo, and KAGRA (LVK) have confirmed about compact binary coalescence (CBC) detections of binary neutron stars, binary black holes and neutron star-black holes Abbott et al. (2019, 2021a, 2021b). The fourth observing run (O4) 111https://observing.docs.ligo.org/plan has seen more than 200 public alerts of gravitational-wave (GW) triggers 222https://gracedb.ligo.org/superevents/public/O4/.
@@ II Data and model architecture
The second dataset is taken from the LIGO, Virgo, and KAGRA (LVK) third observing run (O3) replay Mock Data Challenge (MDC) Sharma Chaudhary et al. (2024). It consists of simulated CBC signals that were analyzed by online LVK search pipelines before the O4 in order to evaluate their performance. We utilize this dataset solely for testing purposes due to the limited number of simulated signals that the pipelines successfully recover. We also utilize a subset of the O3 replay MDC data for our PE analysis in subsection VII.
@@
As the chirp mass of the signal increases, the error in its pipeline recovered value also increases Ewing et al. (2023). Therefore, following a binning approach adopted by the online PE algorithm bilby-pipe 1.3.0, we split the data into four different bins according to the value of the chirp mass. Table 1 lists the lower and upper bounds of these bins as well as their sizes. In order to train and test our method, we further split the datasets into () train, validation. This produces a total of twelve sets.
@@
The advantage of QR over conventional regression is that it can provide conditional probability distributions rather than point estimates Bassett et al. (1978). QR is also less affected by outliers than conventional methods since its loss function is based on absolute errors rather than squared errors.
@@
In the case of a single NN model predicting a number of quantiles , where , the loss function can be generalized to
@@
In our study, the variables denote the CBC parameters for which we want to compute the quantiles (, , and ), with running over the number of samples in each batch. Our NN QR configuration is illustrated in Fig. 1 and described below following the PyTorch Ansel et al. (2024) library nomenclature Foundation (2025). The architecture consists of a fully connected NN with two (24, 12) hidden layers. We choose LeakyReLU Maas (2013) as the activation function for each hidden layer and between each of them we insert a dropout layer to randomly zero out 25% of the nodes in the training process. We normalize and rescale the input before passing it to the network with a LayerNorm layer. The output of the hidden layers is sorted in ascending order by a SoftSort Blondel et al. (2020). This configuration prevents possible inversions of quantile values in the output. The latter is a set of unconstrained real numbers that is conditioned according to the specific training parameter: chirp mass, , mass ratio, , and total component mass . For the mass ratio we apply a sigmoid function to constrain the output of the sorting layer to the physical range . In the case of chirp mass and total component mass, we take the exponential of the sorting layer output and multiply it by the corresponding recovered parameters. For example, in the case of the chirp mass we have , where indicates the output of the sorted layer.
We train the model to produce different quantiles with the stochastic gradient descent and loop through the whole dataset 100 times (epochs), dividing it into mini-batches of samples. The values of the quantiles are selected such that they provide a denser representation at the lower and upper bounds of the quantile range.
We utilize the AdamW Kingma and Ba (2017) optimizer with
the CosineAnnealingWarmRestarts learning rate scheduler. This ensures that the learning rate, which is related to the magnitude of the gradient descent steps, decreases monotonically across the epochs as a cosine curve. The model weights are selected across epochs to minimize the loss on the validation set.
We select the training hyperparameters by testing multiple configurations of the optimizer, learning rate, and the number of nodes in the hidden layers. For the sake of streamlining the training process, we settle on a common choice of hyperparameters that is unique through all bins while keeping the validation loss to a minimum. The initial learning rates are set to () for and ().
@@
Figure 2 displays the loss curves for the twelve independent models (four chirp mass bins three parameters). As the number of the training epoch increases, both training and validation losses decrease, becoming essentially constant after 50 epochs. The validation loss follows the training loss in all models, indicating that there is no overfitting.
@@
To assess the performance of our models we define two metrics, accuracy and interval width. For each quantile with target accuracy , we build the confidence interval , where is any of the training parameters. If the true (injected) value of the parameter lies within the interval, we call accurate. The fraction of accurate intervals across the dataset defines the accuracy of the models. The interval width is defined as . An optimal model produces intervals with maximum accuracy and smallest width compatible with the uncertainties of the recovered parameter.
@@ III O2 results
Figure 3 shows the performance of the models on the O2 testing dataset. The dashed diagonal lines indicate ideal performance. The model for the mass ratio () slightly underperforms for the A and C bins. We pass the final network output for this model onto a sigmoid function to constrain the predictions between and . However, since the sigmoid function only reaches these boundaries asymptotically, the upper bound of the predicted interval is never reached. This leads to a slight bias for the events with .
@@ IV O3 MDC results
| 0.954 | 70.291 |
Figure 5 and Table 3 show the results of the NN models trained on O2 data on the O3 MDC set. Since the parameter space of the O2 and the O3 MDC datasets are different, we evaluate the models on a reduced O3 dataset (see Sec. II.) The different trends of the parameters in the bins with respect to O2 are likely due to the inclusion of multiple pipelines in the O3 dataset, while training on the output of a single pipeline (GstLAL) in the O2 data. Changes in pipelines and their template banks leading up to O4 also affect injection recovery across different datasets. In spite of this, Fig. 5 and table 3 confirm that the NN models perform similarly on the O3 set as they do on O2.
@@ V Performance on a synthetic dataset
In order to assess the NN performance on the full mass parameter space, we create a synthetic dataset. We choose a log-uniform distribution for between 0.5 and 55 to avoid oversampling high values of the chirp mass and a uniform distribution for between 0.05 and 1. As we are mainly interested in assessing the performance on the chirp mass and the mass ratio, we set the spin components to zero and fix the SNR to 10, 15, 20, and 25. Figure 7 shows the NN interval width with 96% target accuracy. Overall, in each bin the interval width decreases as increases. The width also generally increases with at fixed . These trends can also be seen in Fig. 8 which shows the widths of the chirp mass intervals for different values of the SNR. We choose the SNR metric to assess the performance of the NN because it is a common statistic across all pipelines with a clear definition that does not depend on the each pipeline implementations. The median widths decrease with increasing SNR and other input parameters held fixed. This is more evident for bins B, C, and D.
@@ VII NN model outputs in low-latency PE
We produce full posteriors with bilby 2.2.1 and bilby-pipe 1.3.0 and uniform mass priors with intervals computed by the NN. We also perform PE on the same set with the default low-latency bilby listed in Table 5. We employ the relative binning likelihood Krishna et al. (2023) since the low-latency ROQ method is incompatible with the dynamical chirp mass intervals provided by the NN.
@@
Figure 9 compares the number of likelihood evaluations for Bilby’s default configuration and with the NN priors. The NN priors show a % median reduction in the number of likelihood evaluations. Such a reduction is significant in the low-latency scenario, where PE may still require a few hours time even for well localized BBHs.